OpenStudy (anonymous):

What is the equation of the parabola with a directrix at y = -4 and focus (0, 4).

OpenStudy (anonymous):

(x-0)^2 + 3^2- (-1)^2 =2(3+1)

OpenStudy (campbell_st):

well the standard form I use is (x -h)^2 = 4a(y - k) (h, k) is the focus and a is the focal length. start by finding the value of the focal length a. the distance between the focus and directrix is 2a